# ITEACNet: Inverted Teacher-studEnt seArch Conversation Network

### v2

@@ Abstract
Due to the widespread applications of conversations in human-computer interaction, Conversation Emotion Recognition (CER) has attracted increasing attention from researchers. In real-world scenarios, the emotional states of both participants in a conversation tend to maintain a relatively stable pattern within the local context, and often encountering issues with incomplete data patterns. Focusing on these two key challenges, we propose a novel framework for incomplete multimodal learning in CER, called ”Inverted Teacher-studEnt seArch Conversation Network (ITEACNet).” ITEACNet comprises two novel components: the ”Emotion Context Changing Encoder (ECCE)” and the ”Inverted Teacher-Student framework (ITS).” ECCE considers context changes from both local and global perspectives, while the ITS allows a simple teacher model to learn complete data processing methods, enabling a complex student model to follow the performance of the teacher model using incomplete data. Furthermore, we employ a Neural Architecture Search algorithm to enhance the capabilities of student model , achieving superior model performance. Finally, to align with real-world scenarios, we introduce a novel evaluation method, testing the model’s performance under different missing rate conditions without altering the model weights. We conduct experiments on three benchmark CER datasets, and the results demonstrate that our ITEACNet outperforms existing methods in incomplete multimodal CER.
*Equal contribution
Introduction
@@
Conversation Emotion Recognition (CER) aims to analyze the emotional states of speakers in conversational contexts. It has found extensive applications in areas such as social media analysis and human-computer interaction (Lian et al. 2023). Recent studies have achieved promising results in CER. However, there are still some limitations that need to be addressed.
@@
First, unlike sentence-level emotion understanding tasks, conversations feature longer sequences and important context information. Long-sequence conversations often entail complex context changes, the model need to carefully consider these changes to better understand conversation data. Second, real-world scenarios introduce various challenges that impact the collection of multimodal data. Factors such as background noise in speech, obstacles in images, and recognition errors in text may result in the omission of modal information, introducing uncertainty to the completeness of the data. Consequently, the model must exhibit robustness in addressing incomplete data.
@@
In this paper, we propose a novel framework for incomplete learning in CER, called ”Inverted Teacher-studEnt seArch Conversation Network (ITEACNet)”. Fig 1 illustrates the overall structure of our method. To better understand the context changes in conversations, we design a novel encoding method that enables the model to simultaneously consider them from both local and global perspectives, called Emotion Context Change Encoding (ECCE). To enhance the model’s ability to handle missing data, we propose a novel training framework, called Inverted Teacher-Student (ITE), which allows the model to utilize computationally intensive techniques, such as Neural Architecture Search (NAS) (Liu, Simonyan, and Yang 2019), to more fully utilize undamaged data to impute missing data.
@@
To verify the effectiveness of our method, we conduct experiments on three benchmark CER datasets. However, due to the uncertain nature of data incompleteness in real-world scenarios, existing evaluation methods, which involve training model weights separately for testing under various missing rate scenarios (Sun et al. 2023; Lian et al. 2023; Peng et al. 2020), fail to authentically reflect the model’s robustness. Consequently, we propose a novel evaluation method to simulate more realistic missing scenarios. Specifically, we fix the model weights and assess its performance across different missing rates, providing a more accurate measure of the model’s robustness.
@@
Through quantitative and qualitative analysis, we demonstrate that our ITEACNet outperforms currently advanced approaches. The main contributions of this paper are summarized as follows:
@@
  • We design a novel encoding method, called Emotion Context Change Encoding, that enables the model to simultaneously consider the context changes from both local and global perspectives.
@@
  • We propose a novel Inverted Teacher-Student training method as the foundational framework for incomplete data, and propose the utilizing of NAS to enhance the imputation of the student model.
@@
  • Unlike existing works that train individual models for specific missing rates, we propose a novel evaluation approach, namely assessing the average performance of a unified model across all missing rates. This evaluation method is closer to the real-world scenario and fills the gap in current research.
@@
  • Experimental results on three benchmark datasets demonstrate that our method, ITEACNet, is superior to existing state-of-the-art approaches in CER.
Figure 1: The overall structure of Inverted Teacher-studEnt seArch Conversation Network (ITEACNet) with the trimodal setting. This model focuses on the Conversation Emotion Recognition (CER) task based on incomplete data.
Figure 2: Encoder Layer calculation.
Related Work
@@
Learning from incomplete multimodal data is a fundamental research area in machine learning. One simple approach involves data imputation, followed by utilizing existing classification methods on the imputed data. Adding zero vectors or using the mean values of the same modality within a class as missing data are the most straightforward methods (Parthasarathy and Sundaram 2020; Zhang et al. 2020, 2022). However, the performance of such approaches inevitably reaches an upper limit, as they cannot leverage deeper multimodal correlations to improve downstream task performance. Consequently, some studies start leveraging the correlations between different modalities to optimize this process, manifesting as a low-rank data matrix resulting from high correlation (Yang et al. 2017; Liang et al. 2019). Many works, therefore, project the data into a common space using low-rankness (Fan et al. 2017; Liang et al. 2019). Nevertheless, for large data, these methods often incur high computational costs.
@@
With the advancement of deep learning, deep models have achieved breakthroughs in the generation domain, and numerous deep learning-based imputation methods have demonstrated superior capabilities. For instance, Duan et al. (Duan et al. 2014) utilize autoencoders to estimate missing data, achieving commendable results in incomplete multimodal learning. Tran et al.(Tran et al. 2017) further propose the Cascaded Residual Autoencoder (CRA) to enhance the imputation capability of autoencoders. Additionally, Zhao et al. (Zhao, Li, and Jin 2021) combine CRA with cycle-consistency loss for cross-modal imputation, outperforming existing methods. Furthermore, Zuo et al. (Zuo et al. 2023) and Sun et al. (Sun et al. 2023) enhance the imputation ability of autoencoders for missing data from both unimodal and multimodal perspectives, or both high- and low-level features, achieving outstanding performance. Meanwhile, Lian et al. (Lian et al. 2023) delve deeper into considering temporal and speaker information in conversations, achieving state-of-the-art performance in CER.
Certainly, besides imputation, there exist alternative strategies for handling incomplete data. For instance, some approaches partition incomplete data into multiple subsets with complete modalities to model them independently (Yuan et al. 2012; Li et al. 2018). Simultaneously, some works enforce low-dimensional representations for different modalities to ensure high correlation within the same sample (Hotelling 1992; Andrew et al. 2013). However, this work mainly focuses on interpolation methods, so it will not be introduced much.
Method
Input Format
@@
Inverted Teacher-Student
Given that current imputation strategies for missing data mainly focus on the vector distance between complete and incomplete data, dealing with incomplete data introduces an additional imputation task alongside the recognition task for the model. This necessitates more complex computations compared to tasks involving only recognition. Therefore, we design a Inverted Teacher-Student training framework, where two different models learn two types of data. The teacher model requires sufficient recognition capability with complete data, while the student model employs a more complex structure to learn how to process incomplete data.
Teacher-Student training framework is designed to distill knowledge from a complex teacher model into a simpler student model, Inverted Teacher-Student serves as a framework allowing a complex student to emulate the performance of a simple teacher. Our framework is illustrated in Figure 1.
We define complete multimodal features as $\{X_{t},X_{a},X_{v}\}$ and incomplete multimodal data as $\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\}$. We first compute their Emotion Context Change Encoding (ECCE):
$\{E_{t},E_{a},E_{v}\}=\mathrm{ECCEEncoder}(\{X_{t},X_{a},X_{v}\})$ (1)
$\{E_{t}^{s},E_{a}^{s},E_{v}^{s}\}=\mathrm{ECCEEncoder_{s}}(\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\})$ (2)
where $E_{m}\in\mathbb{R}^{L\times L}$. Subsequently, they pass through multiple Encoder Layers in their respective models, generating hidden features under the guidance of ECCE:
@@
In the end, the final layer of hidden features $H_{1}$ and $H_{1}^{s}$ is employed by the classifier to predict emotion labels. Simultaneously, the masked portions of the last three layers of hidden features from the Student Model, $\{H_{1}^{s},H_{2}^{s},H_{3}^{s}\}$, are computed for imputation loss, aiming to closely approximate the vector of the Teacher Model, $\{H_{1},H_{2},H_{3}\}$.
Emotion Context Changing Encoder
@@
As depicted in Figure 3, the emotional state of speakers in a conversation tends to maintain a stable pattern locally, which evolves with context changes. To enhance model’s perception of these dynamics, we design a novel encoding method, Emotion Context Correlation Encoding. The encoder consists of two specific steps: encoding local emotion context and encoding global context changing.
Encoding local emotion context
@@
Encoding global context changing
@@
$F_{\text{global}}^{m}$ represents a replaceable function used for processing context features. After calculating for each utterance pair, we obtain the context changing feature $E_{m}\in\mathbb{R}^{L\times L\times d_{e}}$.
Algorithm 1 Pseudo code of ECCE Encoder.
Input: $X_{m}\in\mathbb{R}^{L\times d_{m}}$, $w$, $dist_{m}$
@@
$F_{\text{local}}^{m}$ = Conv1d($d_{m}$, $d_{z}$, $w$, $pool=w//2$, $stride=1$)
$F_{\text{global}}^{m}$ = Mean()
@@
Algorithm 1 presents the pseudocode for the Emotion Context Changing Encoder. In our experiments, we use a simple 1D convolution operation as $F_{\text{local}}$ and a straightforward average pooling operation as $F_{\text{global}}$.
ITEACNet
To enhance the model’s perception of emotions, we leverage ECCE to guide the modeling of single-modal sequence information. Additionally, to improve the model’s imputation capability for incomplete data, we propose the use of NAS to enhance student. In this section, we will provide a detailed explanation of these processes. As illustrated in Figure 2, the Encoder Layer and ITEAC Encoder subgraphs depict the model we designed, while the Token Mixer subgraph illustrates the guidance process of ECCE.
Encoder Layer
As shown in Figure 2, multimodal features undergo encoding in their respective ITEAC Encoder at each Encoder Layer. Subsequently, the Cross Attention module merges various modal information into text features. The fused text features will serve as the latent vector $H$ in Figure 1. Subsequently, the three modal data will continue to be fed into the next Encoder Layer. The calculation process of Cross Attention can be described as follows:
@@
ITEAC Encoder
In the design of the ITEAC Encoder, we retain the overall framework of Metaformer (Yu et al. 2022b) and conduct a search for the Token Mixer (Yang et al. 2022). Unlike previous workes, to maximize the model’s encoding capacity, we utilize the Router module (Yu et al. 2022a) to extend the architecture search process to input-dependent. Different inputs can choose different architectures as needed, rather than opting for a single architecture for all inputs. Given the candidate module set $\mathbb{O}$, where $\mathbb{|O|}$ represents the number of modules, the search process through the Router can be described as follows:
@@
While there have been numerous excellent designs for Token Mixers recently, our focus is on the introduction of NAS methods rather than algorithm design. Therefore, we select 6 simple and lightweight operations as candidates. Please refer to the appendix for a detailed description of them.
In our experiments, when the Router is a linear layer and the Token Mixer comprises the selected six candidates, we denote the ITEAC Encoder as NAS. Conversely, when the Router is an Identity operation and the Token Mixer includes only Self-Attention, we denote it as Transformer.
Guidance of ECCE
The role of the Token Mixer is to interact with information from different utterances in conversation. A category of mixers, represented by attention operation, typically generates a Mix Map, which is normalized and used as weights for fusing utterance feature. The emotion context change information we designed serves to adjust the Mix Map, thereby guiding the interaction of sequential information (Ying et al. 2021). The specific calculation process can be described as follows:

### v3

# ITEACH-Net: Inverted Teacher-studEnt seArCH Network for Emotion Recognition in Conversation

@@ Abstract
Due to the widespread applications of conversations in human-computer interaction, Emotion Recognition in Conversation (ERC) has attracted increasing attention from researchers. Although many studies have achieved notable results, there remain two critical challenges that hinder the development of ERC. Firstly, there is a lack of exploration into mining deeper insights from the data itself for conversational emotion tasks. Secondly, the systems exhibit vulnerability to random modality feature missing, which is a common occurrence in realistic settings. Focusing on these two key challenges, we propose a novel framework for incomplete multimodal learning in ERC, called ”Inverted Teacher-studEnt seArCH Network (ITEACH-Net).” ITEACH-Net comprises two novel components: the Emotion Context Changing Encoder (ECCE) and the Inverted Teacher-Student (ITS) framework. Specifically, leveraging the tendency for emotional states to exhibit local stability within conversational contexts, ECCE captures these patterns and further perceives their evolution over time. Recognizing the varying challenges of handling incomplete versus complete data, ITS employs a teacher-student framework to decouple the respective computations. Subsequently, through Neural Architecture Search, the student model develops enhanced computational capabilities for handling incomplete data compared to the teacher model. During testing, we design a novel evaluation method, testing the model’s performance under different missing rate conditions without altering the model weights. We conduct experiments on three benchmark ERC datasets, and the results demonstrate that our ITEACH-Net outperforms existing methods in incomplete multimodal ERC. We believe ITEACH-Net can inspire relevant research on the intrinsic nature of emotions within conversation scenarios and pave a more robust route for incomplete learning techniques. Codes will be made available.
keywords:
emotion recognition in conversation , incomplete multimodal learning , emotion context changing ,
inverted teacher-student , neural architecture search
journal: Artificial Intelligence
[…]
[label1]organization=School of Artificial Intelligence, University of Chinese Academy of Sciences,city=Beijing, country=China
[…]
[label2]organization=Institute of Automation, Chinese Academy of Sciences,city=Beijing, country=China
[…]
[label3]organization=University of Science and Technology of China,city=Anhui, country=China
[…]
[label4]organization=Peking University,city=Beijing, country=China
[…]
[label5]organization=Department of Automation,Tsinghua University,city=Beijing, country=China
1 Introduction
@@
Emotion Recognition in Conversation (ERC) aims to analyze the emotional states of speakers in conversational contexts. It has found extensive applications in areas such as social media analysis and human-computer interaction [1, 2, 3]. Recent studies have achieved promising results in ERC [4, 5, 6]. However, there are still some limitations that need to be addressed.
First, unlike sentence-level emotion understanding tasks, conversations feature longer sequences and important context information. Long-sequence conversations often entail complex context changes, the model need to carefully consider these changes to better understand conversation data. While many works have strengthened the ability of models to process context information [7, 8, 9, 10], these methods are general in conversational tasks. They have not deeply considered the unique patterns of emotional expression in conversations, and there is an inability to leverage such prior knowledge to fully exploit data.
@@
Second, real-world scenarios introduce various challenges that impact the collection of multimodal data. Factors such as background noise in speech, obstacles in images, and recognition errors in text may result in the omission of modal information, introducing uncertainty to the completeness of the data. Consequently, the model must exhibit robustness in addressing incomplete data. Although many works have performed well in incomplete learning [11, 12, 13, 14], they overlook the difference in computational power requirements between models for incomplete data tasks and complete data tasks, failing to endow models with sufficient capability to handle missing data.
Third, the disturbances data encounters are uncertain. Therefore, when evaluating the robustness of models, it is essential to consider their performance in scenarios where the incompleteness of the data dynamically changes. Existing evaluation methods, which involve training model weights separately for testing under various missing rate scenarios [15, 16, 17, 18], fail to authentically reflect the model’s robustness.
@@
In this paper, we propose a novel framework for incomplete learning in ERC, called ”Inverted Teacher-studEnt seArCH Network (ITEACH-Net)”. Figure 1 illustrates the overall structure of our method. To better understand the unique patterns in emotional conversations, as described in Figure 3, we design a novel encoder based on the context changes that enables the model to simultaneously consider them from both local and global perspectives, called Emotion Context Change Encoder (ECCE). To enhance the model’s ability to handle missing data, we propose a novel training framework, called Inverted Teacher-Student (ITS), which allows the student model to utilize computationally intensive techniques, such as Neural Architecture Search (NAS) [19, 20, 21, 22], to acquire knowledge from the teacher trained on complete data, thereby enhancing its capability in processing incomplete data.
@@
To verify the actual robustness of models, we propose a novel evaluation method to simulate dynamic data missing scenarios and conduct experiments on three benchmark ERC datasets. Specifically, we fix the model weights and assess its performance across different missing rates.
@@
Through quantitative and qualitative analysis, we demonstrate that our ITEACH-Net outperforms currently advanced approaches. The main contributions of this paper are summarized as follows:
  • 1.
@@
    We introduce a focus on local stable emotional patterns within conversational data, and design a novel encoding method, called Emotion Context Change Encoder, that enables the model to simultaneously consider them from both local and global perspectives.
  • 2.
@@
    We propose a novel Inverted Teacher-Student training method as the foundational framework for incomplete data, which decouples the computation of complete and incomplete data, and leverages NAS to endow the student model with stronger encoding capabilities, thereby improving its robustness in the face of incomplete data.
  • 3.
@@
    Unlike existing works that train individual models for specific missing rates, we propose a novel evaluation approach, namely assessing the average performance of a unified model across all missing rates. This evaluation method is closer to the dynamic data missing scenario and fills the gap in current research.
  • 4.
@@
    Experimental results on three benchmark datasets demonstrate that our method, ITEACH-Net, is superior to existing state-of-the-art approaches in ERC.
Figure 1: The overall structure of Inverted Teacher-studEnt seArCH Network (ITEACH-Net) with the trimodal setting. The Inverted Teacher-Student framework employs a complex student model to learn from a simple teacher model. The Emotion Context Changing Encoder (ECCE) captures the intricate context information within conversations.
Figure 2: The computational modules within the Teacher Model and the Student Model.
2 Related Work
2.1 Context Modeling in Conversation
In comparison to sentence-level emotion recognition tasks, context information in conversations is extremely crucial [7]. Utilizing temporal modeling methods to consider context information is simple and effective, whether using positional encoding [23] or temporal models [24, 25, 8], both can do this. Based on these methods, many works have further enhanced the modeling of context information. Yang et al. [26] introduce context information by embedding previous statements between interlocutors, thereby enhancing the emotional representation of the current utterance. Liu et al. [27] use a BLSTM [28] layer to learn long-term dependencies and utterance-level context information, and a multi-head self-attention layer to make the model focus on the features most related to the emotions. Song et al. [9] innovatively introduce the concept of global emotional atmosphere, using the most dominant emotion in the conversation as guiding information, and use graph neural networks [29] and GRUs [30] in a concatenated framework to consider local and long-term context information. However, these context modeling methods are universally applicable across various conversational tasks, they do not delve into the unique characteristics inherent in conversational emotional data.
Further, Liu et al. [10] consider context modeling from the perspectives of emotional inertia and contagion, designing two independent modules to capture identity-based global context dependencies and extract speaker- and temporal-aware local context information. Tu et al. [31] enhance the model’s ability to capture emotional information by starting from the dependency relationship between commonsense knowledge and context. Peng et al. [32] explore the interrelationship between intention, action, and emotion, while Ai et al. [33] use graph neural networks to model the dependencies between speakers and events in conversations. These methods, although considering the characteristics of emotional information in conversational tasks, primarily focus on the association with external knowledge and lack an in-depth exploration of the intrinsic properties of the data itself.
2.2 Incomplete Multimodal Learning
@@
Learning from incomplete multimodal data is a fundamental research area in machine learning. One simple approach involves data imputation, followed by utilizing existing classification methods on the imputed data. Adding zero vectors or using the mean values of the same modality within a class as missing data are the most straightforward methods [34, 35, 36]. However, the performance of such approaches inevitably reaches an upper limit, as they cannot leverage deeper multimodal correlations to improve downstream task performance. Consequently, some studies start leveraging the correlations between different modalities to optimize this process, manifesting as a low-rank data matrix resulting from high correlation [37, 38]. Many works, therefore, project the data into a common space using low-rankness [39, 38]. These methods have shown promising results in incomplete multimodal learning.
@@
However, with the advancement of deep learning, methods based on deep learning have significantly surpassed the aforementioned approaches. For instance, Duan et al. [40] utilize autoencoders to estimate missing data, achieving commendable results in incomplete multimodal learning. Tran et al.[41] further propose the Cascaded Residual Autoencoder (CRA) to enhance the imputation capability of autoencoders. Additionally, Zhao et al. [11] combine CRA with cycle-consistency loss for cross-modal imputation, outperforming existing methods. Furthermore, Zuo et al. [17] and Sun et al. [14] enhance the imputation ability of autoencoders for missing data from both unimodal and multimodal perspectives, or both high- and low-level features, achieving outstanding performance. Meanwhile, Lian et al. [13] delve deeper into considering temporal and speaker information in conversations, achieving state-of-the-art performance in ERC. Although these works have shown excellent performance in handling incomplete data, using the same network structure to model both complete and incomplete data means that the model lacks additional capabilities to strengthen learning from incomplete data.
3 Method
3.1 Input Format
@@
3.2 Inverted Teacher-Student
To ensure that the model achieves performance akin to that with complete data when processing incomplete information, existing approaches primarily focus on encoding the incomplete data into deep representations to approximate those derived from complete data. This introduces additional computational burden to the model in a multitask format, necessitating greater computational capacity than models operating on complete data. Therefore, we introduce the Teacher-Student framework to decouple these two computations, where the teacher model is trained on complete data, and the student model focuses on utilizing incomplete data to emulate the teacher model’s performance. Furthermore, we design an Inverted Teacher-Student framework, providing the student model with a more complex structure than the teacher, enabling it to possess sufficient computational capacity to follow the teacher’s performance.
Teacher-Student training framework is designed to distill knowledge from a complex teacher model into a simpler student model [42, 43, 44], Inverted Teacher-Student serves as a framework allowing a complex student to emulate the performance of a simple teacher. Our framework is illustrated in Figure 1.
We define complete multimodal features as $\{X_{t},X_{a},X_{v}\}$ and incomplete multimodal data as $\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\}$. We first capture their emotion context changing information:
$\{E_{t},E_{a},E_{v}\}=\mathrm{ECCE}(\{X_{t},X_{a},X_{v}\})$ (1)
$\{E_{t}^{s},E_{a}^{s},E_{v}^{s}\}=\mathrm{ECCE_{s}}(\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\})$ (2)
where $E_{m}\in\mathbb{R}^{L\times L}$. Subsequently, they pass through multiple Encoder Layers in their respective models, generating hidden features under the guidance of $E_{m}$:
@@ 3.2 Inverted Teacher-Student
In the end, the final layer of hidden features $H_{1}$ and $H_{1}^{s}$ is employed by the classifier to predict emotion labels. Simultaneously, the masked portions of the last three layers of hidden features from the Student Model, $\{H_{1}^{s},H_{2}^{s},H_{3}^{s}\}$, are computed for distance loss, aiming to acquire the knowledge from the Teacher Model, $\{H_{1},H_{2},H_{3}\}$.
3.3 Emotion Context Changing Encoder
@@
As depicted in Figure 3, the emotional state of speakers in a conversation tends to maintain a stable pattern locally, which evolves with context changes. To enhance model’s perception of these dynamics, we design a novel encoding method, Emotion Context Changing Encoder (ECCE). The encoder consists of two specific steps: encoding local emotion context and encoding global context changing.
3.3.1 Encoding local emotion context
@@
3.3.2 Encoding global context changing
@@
$F_{\text{global}}^{m}$ represents a replaceable function used for processing context features. During computation, the maximum allowable distance $dist_{m}$ between $i$ and $j$ can be set to mitigate excessive computational complexity. After calculating for each utterance pair, we obtain the context changing feature $E_{m}\in\mathbb{R}^{L\times L\times d_{e}}$.
Algorithm 1 Pseudo code of ECCE.
Input: $X_{m}\in\mathbb{R}^{L\times d_{m}}$, $w$, $dist_{m}$, $F_{\text{local}}^{m}$, $F_{\text{global}}^{m}$
@@
Algorithm 1 presents the pseudocode for the ECCE. In our experiments, we use a 1D convolution operation as $F_{\text{local}}$ and a straightforward average pooling operation as $F_{\text{global}}$.
3.4 Teacher-Student Model
To enhance the model’s perception of complex contexts in conversations, we leverage emotion context changing information $E_{m}$ to guide the modeling of single-modal sequence information. Additionally, to improve the model’s robustness for incomplete data, we propose the use of NAS to enhance the mimetic capabilities of the student model. In this section, we will provide a detailed explanation of these processes. As illustrated in Figure 2, the Encoder Layer and ITEACH Encoder subgraphs depict the model we designed, while the Token Mixer subgraph illustrates the guidance process of $E_{m}$.
3.4.1 Encoder Layer
As shown in Figure 2, multimodal features undergo encoding in their respective ITEACH Encoder at each Encoder Layer. Subsequently, the Cross Attention module merges various modal information into text features. The fused text features will serve as the latent vector $H$ in Figure 1. Subsequently, the three modal data will continue to be fed into the next Encoder Layer. The calculation process of Cross Attention can be described as follows:
@@ 3.4.1 Encoder Layer
3.4.2 ITEACH Encoder
In the design of the ITEACH Encoder, we retain the overall framework of Metaformer [45] and conduct a search for the Token Mixer [20]. Unlike previous workes, to maximize the model’s encoding capacity, we utilize the Router module [46] to extend the architecture search process to input-dependent. As shown in Figure 4, different inputs can choose different architectures as needed, rather than opting for a single architecture for all inputs. Given the candidate module set $\mathbb{O}$, where $\mathbb{|O|}$ represents the number of modules, the search process through the Router can be described as follows:
@@ 3.4.2 ITEACH Encoder
While there have been numerous excellent designs for Token Mixers recently, our focus is on the introduction of NAS methods rather than algorithm design. Therefore, we select 4 simple and lightweight operations as candidates. Please refer to the appendix for a detailed description of them.
In our experiments, when the Router is a linear layer and the Token Mixer comprises the selected four candidates, we denote the ITEACH Encoder as NAS. Conversely, when the Router is an Identity operation and the Token Mixer includes only Self-Attention, we denote it as Transformer.
Figure 4: When there are three Token Mixers, the Router’s search process involves different utterance features selecting different operations. The parameters for Token Mixer i are shared.
3.4.3 Guidance of ECCE
The role of the Token Mixer is to interact with information from different utterances in conversation. A category of mixers, represented by attention operation, typically generates a Mix Map, which is normalized and used as weights for fusing utterance feature. The $E_{m}$ serves to adjust the Mix Map, thereby guiding the interaction of sequential information [47]. The specific calculation process can be described as follows: